4. [6] The table below gives 210Pb activities in a marine sediment core.


Depth, L(cm)

1.0

1.5

2.0

2.5

3.0

3.5

4.0

4.5

5.0

210Pb activity, AL(Bq kg-1)

247.0

199.0

160.0

123.0

101.0

78.0

62.0

46.0

38.0


Assume the following equation applies


Where AL is the activity at depth L (Bq kg-1), A0 is the activity at zero depth (= 0 cm), l is the decay constant (0.031 yr-1 for 210Pb), L is the sediment depth (cm) and S is the sediment accumulation rate (cm yr-1).


(a) Transform the above equation into straight line form.


(b) By plotting the data estimate the value of A0 and S.


1
Expert's answer
2022-02-23T09:23:46-0500

Let t be the age of a layer. Therefore, "L = tS."

The activity is "A_L = A_0e^{-tl}."

So "A_L = A_0e^{-\frac{L}{S}l}." It is an exponential curve. If we take logarithm, we'll get

"\ln A_L = \ln A_0 - \frac{L}{S}l = \ln A_0 - L \cdot\dfrac{l}{S}." It is a linear function of L.

From a linear plot we get "\ln A_0 = 5.9 \; \Rightarrow A_0 = e^{5.9} = 365" Bq kg-1 ,"\frac{l}{S} = 0.475 \Rightarrow S = 0.065" cm/y.